OpenStudy (anonymous):

In a family, each boy has as many sisters as brothers, but each sister has only half as many sisters as brothers. How many children are there in the family?

OpenStudy (goformit100):

ok

OpenStudy (anonymous):

@Decart show me the solution please :)

OpenStudy (anonymous):

Ah ok I got it by fingers. But is there any mathematecal equation to solve this kind of problem? Its like a riddle.

OpenStudy (anonymous):

boys = x girls = y

OpenStudy (campbell_st):

there are 7 children, 4 boys and 3 girls...

OpenStudy (anonymous):

@campbell_st can you show me your solution?

OpenStudy (campbell_st):

plenty of options...

OpenStudy (cwrw238):

i find this quite tricky..

OpenStudy (anonymous):

two equations two unknowns

OpenStudy (campbell_st):

if you rewrite Decarts equations you have y = -1 and 0.5y - x = 1

OpenStudy (anonymous):

oh wait. im still trying to figure it out. its tricky :/

OpenStudy (campbell_st):

for me it was trial and error knowing that there needed to be an odd number of children I knew 3 doesn't work, 5 works for the boy having equal sisters and brothers but doesn't work for the girl, so try 7 works

OpenStudy (cwrw238):

- thats probably the best way i've got a headache trying to create the equations

OpenStudy (anonymous):

but we can get them

OpenStudy (cwrw238):

maybe i need another coffee - lol

OpenStudy (anonymous):

@cwrw238 haha, my brother too. until he left me here alone :/

OpenStudy (anonymous):

brother = x+y sister=x+.5y

OpenStudy (anonymous):

.5y when we took one away

OpenStudy (anonymous):

sister = 2x+y

OpenStudy (anonymous):

brother is actually x+1

OpenStudy (anonymous):

x+1brother has x-1+y x-1 being himself

OpenStudy (cwrw238):

1 + 2x = x + y where x = girls, y = boys is one equation

OpenStudy (anonymous):

x+1+y=y-1+2x

OpenStudy (cwrw238):

that gives x + 1 = 1 + 2x giving x=0

OpenStudy (anonymous):

i formulated an equation XD x - 1 = y y - 1 = x/2

OpenStudy (cwrw238):

x - 1 = 0.5y is another so u have 1 + 2x = x + y x -1 = 0.5y is that right??

OpenStudy (anonymous):

@cwrw238 i'll try to solve yours & you try mine :)

OpenStudy (cwrw238):

yes that gives x = 3 y=4

OpenStudy (cwrw238):

oh ok

OpenStudy (anonymous):

awesome

OpenStudy (cwrw238):

x - 1 = y y - 1 = 0.5x add x - 2 = 0.5x x = 4 y = 4-1 = 3

OpenStudy (cwrw238):

in your equation x = boys , y = girls right? yours come out right too

OpenStudy (anonymous):

ya and yours is x = girls and y = boys,right?

OpenStudy (cwrw238):

yea well done

OpenStudy (anonymous):

thank you @cwrw238 :)